SpaceRef

SpaceRef


SOHO Daily Meeting Minutes for Monday 23 June 2003

Status Report From: Goddard Space Flight Center
Posted: Monday, June 23, 2003

ANNOUNCEMENTS

* Current RHESSI default target is AR10386. When selecting this target to maximize coordination w/others, use campaign number 6850.

* EIT Shutterless Segment #2 is tomorrow. No NRT will be available from 19-21:40 UT.

* GITT scans for the EOF and EAF are scheduled for Thursday 26 June at 9:30am.

FOT REPORT

Spacecraft Status: Nominal, HGA Position Monitoring Disabled, HGA Nominal Control Disabled

Spacecraft Anomalies: 172/1357 CEPAC ESU Data Request Error; recovered autonomously after 4 minutes. 174/0357 Single DMA failure over 1 format occurred during heavy NRT commanding by LASCO.

Accomplished Activities: DOY 171: VIRGO, SWAN, MDI, RSL, New tracking star, mag = 7.45 OCD #1795: Lower APMM Thermal Settings OCD #1796: HGA Z-Axis Movement at Low APME Temps OCD #1797: Stabilize APMM Temperatures DOY 172: CELIAS OCD #1798: Z-Axis High Speed Movement with APME-A, Stable Motor Temps DOY 173: CELIAS

Planned Activities: CELIAS, SWAN, NONCOHO_03174

Upcoming Operations: None

Ground Anomalies:

171/0435 Delayed carrier operations on D66 due to problems with their transmitter. Lost 11 minutes of command ability.

171/1403 Received degraded TLM from D27 from 1303 to 2120 due to bandwidth overflow at JPL. Data should be recoverable from the CDR.

171/1439 Commands were being corrupted at D16 due to problems with their command modulation settings. Pass activities delayed 36 minutes.

171/1700 DPS POCC products could not be delivered due to a file system problem. Delivery of POCC products delayed. No real-time operational impacts.

172/0403 SH4WS2 terminal hung, cause unknown. Operations temporarily moved over to String 5. SH4WS2 down from 0403 until 0438.

172/1405 Received degraded TLM from D16 from 1405 until 0130 due to a line overflow at JPL, cause unknown. MDI high rate data collected in VC3 vice VC2. All data recoverable from the CDR.

172/1644 Lost a total of 2.5 minutes of telemetry due to problems with auto-track at D16. Lost data is unrecoverable.

173/0440 D66 antennna RED due to a MPA failure. Lost 530 minutes of TLM and CMD ability. Data recoverable from the SSR and from D54 supplemental support. MDI continuous campaign disrupted from 0440 to 1239.

173/1412 Received degraded TLM from D16 from 1412 to 1800 due to a line overflow at JPL. Data should be recoverable from the CDR.

174/0755 Received degraded TLM from D66 from 0755 until 1026, cause unknown. Poor data quality confirmed by JPL. Lost data should be recoverbale from the CDR.

SOLAR STATUS

06/20 DOY 171 - 06/23 DOY 174

EIT daily On-Line Solar Status and observations: http://umbra.nascom.nasa.gov/eit/plan/log_form20030623.html

LASCO daily On-Line Solar Status and observations: http://lasco-www.nrl.navy.mil/observations/Todays_obs.html

Web Page for Planning: http://www.bbso.njit.edu/arm

Magnetic Maps and Heliospheric Forecasts: http://www.lmsal.com/forecast/

INSTRUMENT STATUS

CDS: Nominal. Engineering studies, sunspot oscillation studies. Tomorrow: TBD and safing.

UVCS: Nominal. JOP155 and synoptics then safing.

LASCO: Nominal. C2/C3 synoptics, JOP155. Safing tomorrow.

EIT: Nominal. 195 CME watch and synoptics, JOP155. Tomorrow: shutterless then safing.

MDI: Nominal. Full disk magnetograms and dopplergrams. Tomorrow: EIT shutterless support then safing.

TRACE: Nominal. Flarewatch. Tomorrow: EIT shutterless support.

AOB

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.